OpenStudy (anonymous):

the tens digit of a two digit number is 3 less than its unit digit and the number is 4 times the sum of the digits. what is the number?

OpenStudy (anonymous):

let "t" be the tens digit let "u" be the unit digit let "x" be the number t=u-3 x=4(t+u)

OpenStudy (anonymous):

now for some reasoning x must be a multiple of 4, thus meaning that u must be even the lowest number u can be is 4, any lower and the problem wouldn't make sense

OpenStudy (anonymous):

multiple of 4 means u must be even

OpenStudy (anonymous):

thanks anyway. the answer is 36. :)

OpenStudy (anonymous):

right